3. If the coefficient of friction between tires and roadway is 0.40, what is the minimum radius at which a car can turn on a horizontal road when its speed is 43.75 ft/s?
1
Expert's answer
2020-11-17T13:03:31-0500

The force of friction defines the centripetal acceleration the car can withstand:


"\mu mg=ma_c,\\\space\\
\mu mg=m\frac{v^2}{R},\\\space\\
R=\frac{v^2}{\mu g}=149\text{ ft}."